Bond Valuation with Semiannual Payments

Renfro Rentals has issued bonds that have a 12% coupon rate, payable semiannually. The bonds mature in 11 years, have a face value of $1,000, and a yield to maturity of 8%. What is the price of the bonds? Round your answer to the nearest cent.

$______?
Answer #1

As Semiannual Payments, therefore

Coupon Rate = 12%/2 = 6%

Coupon = $1000*6% = $60

Payments = 11*2 = 22

YTM = 8%/2 = 4%

Bond Price = Interest * PVIFA(4%,22) + Par Value * PVIF(4%,22)

  = 60 * 14.45111 + 1000 * 0.42196

Bond Price = $1,289.027 or $1,289.03